# Adversarial Reasoning for Repair Based on Inferred Program Intent

### v1

@@ 4.3. RQ2: Adversarial Intents Effectiveness
Figure 4. Adversarial degree between the inferred intents
First, we quantify the adversarial score between different inferred intents. As shown in Figure 4, the majority of adversarial scores fall within the range of 28.6% to 60.0%. The average adversarial score ranges from 43.8% to 46.5%. This data confirms the adversarial nature between different intents.
Second, we measure the extent to which the generated intents align with the expected program intent. We conduct a quality analysis by manually comparing the inferred intents for 300 bugs against the ground-truth patches. Table 4 presents the analysis, showing how well the inferred intents match the ground truth expected intent. Our analysis reveals that in 18.3% of cases (as shown in the last row), none of the inferred program intents are correct. However, it also highlights the significant benefits of using multiple intents. If only a single inferred intent were used, the best single-intent match (first intent) achieves an accuracy of 62.0% (sum of 25.0%, 12.7%, 18.6%, and 5.7%). In contrast, when leveraging multiple intents, at least one correct match occurs in 81.7% of cases.
Table 5. Adversarial Intent Effectiveness
@@
Fault localization performance improves as adversarial intents encourage exploration at both function and statement levels, extending beyond the provided class. First, adversarial reasoning examines a broader and more varied search space for faulty code, not merely by generating more candidate faulty statements. Our experiment shows that simply increasing the number of candidates does not necessarily improve the final result. Second, adversarial reasoning guides the LLM to locate buggy statements outside functions, such as variable definitions, constructors, and other classes. For example, the bug Math-104, involving a buggy variable definition, was identified only through adversarial prompting and was missed in the initial response, which was limited to the function level. Similarly, the bug Math-22 is fixed as AdverIntent-Agent generates patches across two classes, rather than confining itself to the provided class.
Patch generation performance improves as adversarial intents increase the likelihood of the search space containing a correct patch. We observe that successful patches benefit from prompts that are both adversarial and precise. The number of correctly repaired bugs increased from 36 to 43 by incorporating adversarial intents. Adversarial intents explore a diverse range of expected program behaviors, avoiding repeated flawed attempts and effectively pinpointing root causes with accuracy. This approach is a key difference from prior conversational methods, which focus on limited diversity.
We also observe that adversarial intents greatly benefit from an agent-driven search and decision loop When AdverIntent-Agent creates prompts that encourage the LLM to consider adversarial reasoning, the LLM occasionally requests additional information, such as a new class file. The agent then provides this extra context, enabling exploration of different patch possibilities and helping avoid dead-end refusal responses. Without agent-like interaction, the bug repair loop is more likely to halt prematurely due to adversarial reasoning.
@@ 4.4. RQ3: Tests and Overfitting
Table 6. Generated Test cases and Overfitting Alleviation
Test Cases Likely-overfitting Detected
Initial Low-Confidence Rank Threshold Valid
Defects4J 2.0 25050 28.4% 70% 3916 12
HumanEval-Java 4920 23.8% 70% 1666 8
Table 6 shows the number of tests generated by Agenttest and the number of bugs with overfitting patches discarded by adversarial tests, where overfitting is associated with the original tests. The second column shows the initial number of generated tests, while the third column indicates the number of low-confidence assertions, as identified by evaluating the correctness of assertions in generated patches, and we prioritize tests based on the LLM’s assertion confidence. The final valid number of tests and the number of bugs with discarded overfitting patches are given in the sixth and seventh columns, respectively.
We make the following observations from Table 6. First, AdverIntent-Agent discards overfitting patches for 12 bugs in Defects4J 2.0 and 8 bugs in HumanEval-Java, confirming the effectiveness of adversarial testing in alleviating overfitting. Second, test generation suffers from a low compilation rate (see the third column), where the successful compilation rate varies between complex, real-world programs and simpler ones. We found that many common errors stem from missing dependencies that cause non-compilation, suggesting potential integration with existing test improvement work (test-issta23, ; unit-test-issta24, ; Greybox-issta23, ). Third, using criticism prompts to ask LLMs to double-check assertions and identify potentially incorrect ones is effective, helping us filter out 23.8% to 28.4% of low-confidence tests.
Answer to RQ3: AdverIntent-Agent effectively discards overfitting patches for 12 bugs in Defects4J 2.0 and 8 bugs in HumanEval-Java. Test generation faces a low compilation rate, with variation between complex and simpler programs, largely due to missing dependencies. Yet, using criticism prompts to double-check assertions filters out over 20% of low-confidence tests, improving test reliability.
@@ 4.5. RQ4: Cost of AdverIntent-Agent
Figure 5 in left shows box plot of the token cost across four approaches: ChatRepair, RepairAgent, Cigar, and AdverIntent-Agent. Each box shows the distribution of token costs, with the minimum, first quartile, median, third quartile, and maximum values marked. The figure in right illustrates the cost in cents associated with the median token cost for each approach based on the estimated cost of model GPT-4o. AdverIntent-Agent uses an average of 248K tokens per bug, resulting in a cost of 7.44 cents. We now explain the reason that AdverIntent-Agent has higher costs than RepairAgent and Cigar but lower costs than ChatRepair.
AdverIntent-Agent shows higher costs than RepairAgent due to the realistic fault localization and test generation processes, while AdverIntent-Agent assumes the perfect fault localization is known. AdverIntent-Agent costs are also higher than Cigar’s, due to the prompts of reasoning program intents and test generation, which are key innovations in our approach. However, considering only patch generation, AdverIntent-Agent is more cost-effective than Cigar, as it requests one patch per program intent and root cause, whereas Cigar samples 50 candidates per iteration.
Nonetheless, AdverIntent-Agent remains less costly than ChatRepair, primarily because ChatRepair’s iterative approach includes all historical data in prompts, leading to high token accumulation as iterations reach hundreds. In contrast, AdverIntent-Agent limits itself to three rounds of test and patch refinements, which significantly reduces the total token cost.
@@
Figure 5. Token and Monetary Cost of Repairing Bugs in Defects4J 2.0
@@ 5.1. Automated Program Repair
Code Transformation. Source code transformation to alter program behavior has been the main research focus of many search-based (sofix, ; genpro2009, ; le_goues_genprog_2012, ; Yuan2017ARJAAR, ; Arja-e-Tosem20, ; ssFix, ; long_staged_2015, ; tbar, ; acs, ; martinez2016astor, ; saha2019harnessing, ), constraint-based (semfix, ; Angelixicse16, ; nopol, ; concolic-repair-PLDI21, ; relifix-shinwei-icse15, ; directfix, ), and learning-based repair (tse24-Search-Generate-Modify, ; codit-tse20, ; CoCoNuT, ; DLFix, ; SEQUENCER, ; CURE-icse21, ; jiang2023knod, ; RewardRepair-icse22, ) techniques for a decade. The emphasis lies in constructing syntactically and semantically correct code transformations, taking into account their expressiveness and effectiveness.
Rich Information Beyond Code. Recently, more work considered information beyond code transformation, by including commit message (modit-ase21, ), compiler and test feedback (ye2022selfapr, ; kong2024contrastrepair, ), domain knowledge (tosem24-domain, ; ye2022selfapr, ) and iterative patch refinement (ITER, ). Chen et al. (chen2024teaching, ) improves LMs code generation accuracy by injecting feedback messages generated by the LM itself based on its code execution results.
Conversational-based Repairs. Recent advancements in Large Language Models (LLMs) have paved the way for innovative approaches to automated program repair (APR). ChatRepair(ChatRepair-ISSTA24, ) pioneered a conversational repair method based on LLMs, iteratively generating patches by collecting bug context information and patch execution results. Cigar(hidvégi2024cigar, ) further optimized the repair process by designing prompts and reboot strategies to generate diverse patches while reducing token consumption. ContrastRepair (kong2024contrastrepair, ) takes an LLM-based iterative approach for APR. Their prompt includes a failing test and a similar passing test, where a passing test is either selected from the existing suite or generated using type-aware mutation. CHATDBG (levin2024chatdbg, ) shifted the focus from fully automated repair to a co-pilot for debugging, integrating LLMs with standard debuggers to enhance their capabilities.
Patch Correctness Assessment. Existing methods for assessing overfitting patches typically classify these patches separately and address them in post-processing. Prior work mainly focuses on static code analysis (ye2021ods, ; TianASE20, ; ASE20Wang, ; anti-pattern, ) and test generation and execution analysis (gao2019crash, ; tian2021failtest, ; ICSE18-patchsim, ; issta17-difftgen, ).
@@ 5.2. LLM-based Agent for Code
There is a set of parallel LLM-based agents designed for repairing software issues related to our research (yang2024sweagent, ; zhang2024codeagent, ; autocoderover, ). Yang et al. (yang2024sweagent, ) propose SWE-Agent for repairing GitHub issues. AutoCodeRover (autocoderover, ) combines LLMs with advanced code search techniques to address GitHub issues through program modification or patch generation. Zhang et al. (zhang2024codeagent, ) introduce CodeAgent, a novel LLM-based agent framework that employs external tools for repository-level code generation, integrating five programming tools for information retrieval, code symbol navigation, and code testing. Bouzenia et al. (repairagent, ) propose RepairAgent, which treats the LLM as an agent capable of autonomously planning and executing actions to fix bugs by invoking suitable tools. Our work similarly uses a decision-action-planning loop with multiple iterations and dynamic prompts based on command outputs, with the novel component of reasoning about multiple adversarial intents to explore diverse solutions.
@@ 5.3. Program Reasoning
Program intent reasoning is a critical area of research within software engineering, encompassing various techniques and methodologies aimed at understanding, analyzing, and improving software programs. One notable approach is symbolic execution (varfix, ; sergey-test-equivalence-tosem18, ), which systematically explores program paths to identify potential errors or vulnerabilities. Formal methods provide mathematical techniques for reasoning about program correctness and behavior (SMT, ). SpecRover by Ruan et al. (SpecRover-icse25, ) is a closely related work that extracts code intents via LLMs. This work shares a similar idea with ours by first extracting program intent and then using it to guide patch generation. However, our approach conceptually generates multiple diverse program intents adversarially, maximizing the likelihood that at least one inferred program is correct, whereas SpecRover infers only the most likely program intent.
In the general setting of machine learning for program reasoning, several works propose Chain-of-Thought (yang2023chainofthought, ), Tree-of-Thought (yao2023treeofthoughts, ) and Graph-of-Thought (Graph-of-Thoughts, ) to logical reasoning code generation tasks by breaking them down into understandable intermediate steps, enabling LLMs to handle each step individually. Unlike prior work, our approach reasons about program intent at the function specification level and incorporates multiple adversarial intents to ensure diverse solutions are explored, which is novel to our knowledge.
@@ 6. Threats to Validity
We acknowledge several potential threats to validity: 1) Manual patch validation. The first internal threat arises from the manual validation of believed-correct patches against the reference developer patch. To address this, we carefully examined and discussed each patch, and we make the patches publicly available. 2) Data leakage. Using GPT-4o for AdverIntent-Agent raises data leakage concerns due to its undisclosed training data; however, evaluation on the HumanEval-Java dataset helps strengthen internal validity. 3) Non-deterministic of LLMs. Non-deterministic language model outcomes pose a threat. We mitigate this by evaluating a substantial number of bugs and providing interaction logs for reproducibility and transparency. 4) Limited programming languages. Our experiments are limited to Java projects, which threatens external validity, as results may not generalize to other programming languages or projects. To mitigate this, we include two benchmarks, Defects4J 2.0 and HumanEval-Java, covering 181 diverse projects with various bug types, lengths, and complexities, which strengthens the validity of our findings.

### v2

@@ 4.3. RQ2: Adversarial Intent Effectiveness
Figure 4. Adversarial degree between the inferred intents.
We analyze the effectiveness of adversarial intents based on 300 randomly selected bugs from the Defects4J benchmark. First, we quantify the adversarial score between different inferred intents. As shown in Figure 4, the majority of adversarial scores fall within the range of 28.6% to 60.0%. The average adversarial score ranges from 43.8% to 46.5%. This data confirms the adversarial nature between different intents.
Second, we measure the extent to which the generated intents align with the expected program intent (ground truth intent). Table 4 presents the analysis, showing how well the inferred intents match the expected ground truth intent. As shown in the two last rows of Table 4, our analysis indicates that 81.7% of cases with at least one inferred intent are aligned with the ground-truth intent, while only 18.3% of cases have no correctly inferred program intents.
Specifically, it highlights the significant benefits of using multiple intents. If only a single inferred intent is used, the best single-intent match (the first intent) achieves an accuracy of 62.0% (the sum of 25.0%, 12.7%, 18.6%, and 5.7%). In contrast, when leveraging multiple intents, at least one correct match is found in 81.7% of cases. The second and third inferred intents complement the first intent, accounting for the remaining 19.7% of matches.
Fault localization performance improves as adversarial intents encourage exploration at both function and statement levels, extending beyond the provided class. First, adversarial reasoning examines a broader and more varied search space for faulty code, not merely by generating more candidate faulty statements. Our experiment shows that simply increasing the number of candidates does not necessarily improve the final result. Second, adversarial reasoning guides the LLM to locate buggy statements outside functions, such as variable definitions, constructors, and other classes.
Table 5. Adversarial intent effectiveness.
@@
Patch generation performance improves as adversarial intents increase the likelihood of the search space containing a correct patch. We observe that successful patches benefit from prompts that are both adversarial and precise. The number of correctly repaired bugs increased from 36 to 43 by incorporating adversarial intents. Adversarial intents explore a diverse range of expected program behaviors, avoiding repeated flawed attempts and effectively pinpointing root causes with accuracy.
@@ 4.4. RQ3: Tests and Overfitting
Table 6 shows the number of tests generated by Agenttest and the number of bugs with overfitting patches discarded by adversarial tests, where overfitting is associated with the original tests. The second column shows the initial number of generated tests, while the third column indicates the number of low-confidence assertions, as identified by evaluating the correctness of assertions in generated patches, and we prioritize tests based on the LLM’s assertion confidence. The number of bugs with likely-overfitting patches that are discarded are given in the last column.
We make the following observations from Table 6. First, AdverIntent-Agent discards likely-overfitting patches for 12 bugs in Defects4J 2.0 and 7 bugs in HumanEval-Java, confirming the effectiveness of adversarial testing in alleviating overfitting. Second, using criticism prompts to ask LLMs to double-check assertions and identify potentially incorrect ones is effective, helping us filter out 23.8% to 28.4% of low-confidence tests.
Answer to RQ3: AdverIntent-Agent effectively discards overfitting patches for 12 bugs in Defects4J 2.0 and 7 bugs in HumanEval-Java. Test generation faces a low compilation rate, with variation between complex and simpler programs, largely due to missing dependencies.
@@ 4.5. RQ4: Cost of AdverIntent-Agent
Table 7 compares the average token cost across four approaches: Cigar, RepairAgent, ChatRepair, and AdverIntent-Agent. On average, AdverIntent-Agent uses 438K tokens per bug. Its token usage is higher than both RepairAgent and Cigar but remains lower than ChatRepair.
The main reason for the increased cost compared to RepairAgent is that AdverIntent-Agent performs realistic fault localization and automated test generation for each bug, with test generation representing the primary contributor to the overall token cost, a step that is not included in RepairAgent’s workflow.
Compared to Cigar, AdverIntent-Agent ’s higher token cost results from its use of advanced prompts for reasoning about program intent and generating tests, both of which are key innovations in our approach. However, if we consider only the patch generation phase, AdverIntent-Agent is actually more cost-effective than Cigar because it requests three patches per program intent and root cause, whereas Cigar samples 50 candidate patches per iteration, leading to a much greater token expenditure for Cigar in this stage.
Despite these additional costs, AdverIntent-Agent remains more efficient than ChatRepair. ChatRepair’s iterative approach incorporates all historical information into each prompt, causing token consumption to increase rapidly as iterations accumulate, sometimes reaching into the hundreds. By contrast, AdverIntent-Agent is designed to limit itself to three rounds of test and patch refinements, which manages overall token usage
AdverIntent-Agent achieves a balance between realism and efficiency, using more tokens for enhanced reasoning and test generation than some baselines, but remaining substantially more cost-effective than highly iterative methods like ChatRepair.
Table 6. Generated test cases and overfitting detection.
Test Cases Likely-overfitting
Initial Low-Confidence Rank Threshold
Defects4J 2.0 25050 28.4% 70% 12
HumanEval-Java 4920 23.8% 70% 7
@@ 5.1. Automated Program Repair
Code Transformation. Source code transformation to alter program behavior has been the main research focus of many search-based (Liu and Zhong, 2018; Weimer et al., 2009; Le Goues et al., 2012b; Yuan and Banzhaf, 2018, 2020; Xin and Reiss, 2017b; Long and Rinard, 2015; Liu et al., 2019; Xiong et al., 2017; Martinez and Monperrus, 2016b; Saha et al., 2019), constraint-based (Nguyen et al., 2013; Mechtaev et al., 2016; Xuan et al., 2016; Shariffdeen et al., 2021a; Tan and Roychoudhury, 2015; Mechtaev et al., 2015), and learning-based repair (Liu et al., 2024; Chakraborty et al., 2020; Lutellier et al., 2020; Li et al., 2020; Chen et al., 2019; Jiang et al., 2021, 2023b; Ye et al., 2022a) techniques for a decade. The emphasis lies in constructing syntactically and semantically correct code transformations, taking into account their expressiveness and effectiveness.
Rich Information Beyond Code. Recently, more work considered information beyond code transformation, by including commit message (Chakraborty and Ray, 2021), compiler and test feedback (Ye et al., 2022b; Kong et al., 2024), domain knowledge (Zirak and Hemmati, 2024; Ye et al., 2022b) and iterative patch refinement (Ye and Monperrus, 2024). Chen et al. (Chen et al., 2024) improves LMs code generation accuracy by injecting feedback messages generated by the LM itself based on its code execution results.
Conversational-based Repairs. Recent advancements in Large Language Models (LLMs) have paved the way for innovative approaches to automated program repair (APR). ChatRepair(Xia and Zhang, 2024) pioneered a conversational repair method based on LLMs, iteratively generating patches by collecting bug context information and patch execution results. Cigar(Hidvégi et al., 2024) further optimized the repair process by designing prompts and reboot strategies to generate diverse patches while reducing token consumption. ContrastRepair (Kong et al., 2024) takes an LLM-based iterative approach for APR. Their prompt includes a failing test and a similar passing test, where a passing test is either selected from the existing suite or generated using type-aware mutation. CHATDBG (Levin et al., 2024) shifted the focus from fully automated repair to a co-pilot for debugging, integrating LLMs with standard debuggers to enhance their capabilities.
Patch Correctness Assessment. Existing methods for assessing overfitting patches typically classify these patches separately and address them in post-processing. Prior work mainly focuses on static code analysis (Ye et al., 2021a; Tian et al., 2020; Wang et al., 2020; Tan et al., 2016) and test generation and execution analysis (Gao et al., 2019b; Tian et al., 2022; Xiong et al., 2018; Xin and Reiss, 2017a).
@@ 5.2. LLM-Based Agent for Code
There is a set of parallel LLM-based agents designed for repairing software issues related to our research (Yang et al., 2024; Zhang et al., 2024a, b). Yang et al. (Yang et al., 2024) propose SWE-Agent for repairing GitHub issues. AutoCodeRover (Zhang et al., 2024b) combines LLMs with advanced code search techniques to address GitHub issues through program modification or patch generation. Zhang et al. (Zhang et al., 2024a) introduce CodeAgent, a novel LLM-based agent framework that employs external tools for repository-level code generation, integrating five programming tools for information retrieval, code symbol navigation, and code testing. Bouzenia et al. (Bouzenia et al., 2024) propose RepairAgent, which treats the LLM as an agent capable of autonomously planning and executing actions to fix bugs by invoking suitable tools. Our work similarly uses a decision-action-planning loop with multiple iterations and dynamic prompts based on command outputs, with the novel component of reasoning about multiple adversarial intents to explore diverse solutions.
Table 7. Token cost analysis of AdverIntent-Agent.
Cigar RepairAgent ChatRepair AdverIntent-Agent
Tokens 127K 270K 467K 438K
@@ 5.3. Program Reasoning
Program intent reasoning is a critical area of research within software engineering, encompassing various techniques and methodologies aimed at understanding, analyzing, and improving software programs. One notable approach is symbolic execution (Wong et al., 2021; Mechtaev et al., 2018), which systematically explores program paths to identify potential errors or vulnerabilities. Formal methods provide mathematical techniques for reasoning about program correctness and behavior (Barrett et al., 2009). SpecRover by Ruan et al. (Ruan et al., 2025) is a closely related work that extracts code intents via LLMs. This work shares a similar idea with ours by first extracting program intent and then using it to guide patch generation. However, our approach conceptually generates multiple diverse program intents adversarially, maximizing the likelihood that at least one inferred program is correct, whereas SpecRover infers only the most likely program intent.
In the general setting of machine learning for program reasoning, several works propose Chain-of-Thought (Yang et al., 2023), Tree-of-Thought (Yao et al., 2023) and Graph-of-Thought (Besta et al., 2024) to logical reasoning code generation tasks by breaking them down into understandable intermediate steps, enabling LLMs to handle each step individually. Unlike prior work, our approach reasons about program intent at the function specification level and incorporates multiple adversarial intents to ensure diverse solutions are explored, which is novel to our knowledge.
@@ 6. Threats to Validity
We acknowledge several potential threats to validity: 1) Manual patch validation. The first internal threat arises from the manual validation of believed-correct patches against the reference developer patch. To address this, we carefully examined and discussed each patch, and we make the patches publicly available. 2) Data leakage. Using GPT-4o for AdverIntent-Agent raises data leakage concerns due to its undisclosed training data; however, evaluation on the HumanEval-Java dataset helps strengthen internal validity. 3) Non-deterministic of LLMs. Non-deterministic language model outcomes pose a threat. We mitigate this by evaluating a substantial number of bugs and providing interaction logs for reproducibility and transparency. 4) Limited programming languages. Our experiments are limited to Java projects, which threatens external validity, as results may not generalize to other programming languages or projects. To mitigate this, we include two benchmarks, Defects4J 2.0 and HumanEval-Java, covering 181 diverse projects with various bug types, lengths, and complexities, which strengthens the validity of our findings. 5) Hallucinations and mistakes of LLMs. We use LLMs for test generation based on specific program intent, but the generated oracles can sometimes be incorrect and difficult to detect. LLMs are also used to identify overfitting patches, which may result in further mistakes. We mitigate this risk by manually reviewing the results of LLM-based detection.